Calculate the volume of oxygen at 15°c and 745mmhg which could be obtained by heating 10g of potassium trioxochlorate. ( K =39, Cl=35.5, O=16. Molar mass of gas at STP=22.4dm3
1
Expert's answer
2019-11-22T04:34:39-0500
"2KClO_3 = 2KCl + 3O_2"

"n (KClO_3) = m\/M = 10\/122.5 = 0.082mol"

"n(O_2) = 3\/2 * 0.082 = 0.123mol"

"V(O_2) = n*V_m = 0.123*22.4 = 2.755dm^3"

From the combinated gaseous law:


"V^`(O_2) = {T_2p_1V(O_2) \over T_1p_2}"

"V^`(O_2) = {288.15*760*2.755 \over 273.15*745} = 34.982dm^3 \approx 35dm^3"

Answer: 35dm3